Town M and town N are connected by two roads while town N and town P are connected by five roads. How many different routes are there from town M to town P through town N? State your solution and explain


1
Expert's answer
2021-11-30T04:21:55-0500

There are 2 different roads connecting town M and town N. This shows that there are "k=2" different routes from M to N.

There are 5 different roads connecting town N and town P. Thus, there are "r=5" different routes from N to P.

Therefore, there are "k*r=2*5=10" different routes from town M to town P through town N.